OpenStudy (anonymous):

Solve for x a) log4 (x+4) - log4 (x-2) = 3 b) 4^3x = 42 c) 700(1+0.06/52)^52x = 1275

sam (.sam.):

a) log_4(x+4)-log_4(x-2)=3 log_4((x+4)/(x-2))=3 4^(log_4((x+4)/(x-2)))=4^(3) (x+4)/(x-2)=64 (x+4)/(64)=(x-2) (x+4)/(64)=x-2 Least common denominator: 64 x+4=64x-128 -63x+4=-128 -63x=-132 x=44/21

sam (.sam.):

b) 4^(3x)=42 ln(4^(3x))=ln(42) xln(64)=ln(42) x=ln(42)/ln(64) x=0.8987

sam (.sam.):

c) If your "c" is in the form 700(1+(0.06)/(52))^(5)(2x)=1275 2x=1.811 x=0.9055